OpenStudy (anonymous):

19. Tom has a collection of 30 CDs and Nita has a collection of 18 CDs. Tom is adding 1 CD a month to his collection while Nita is adding 5 CDs a month to her collection. Find the number of months after which they will have the same number of CDs. Show your work.

OpenStudy (cutiecomittee123):

Well first make an equation for each person

OpenStudy (anonymous):

Okay I'd appreciate if you would walk me through it

OpenStudy (cutiecomittee123):

Since Tom has 30 CDs you know that the end of the equation will have a +30

OpenStudy (anonymous):

okay

OpenStudy (cutiecomittee123):

and the same goes for nita the end of her equation would have a +18

OpenStudy (anonymous):

@cutiecomittee123 are you still helping?

OpenStudy (cutiecomittee123):

Sorry I had to go to my own question, but sure Ill try. So now that you have the constants, you are going to need to put in a change. What types of equations are you learning in school right now? like what math are you in?

OpenStudy (cutiecomittee123):

It would help to know that in order to get a basis for what the equation is supposed to be set up as.

OpenStudy (anonymous):

This is my algebra exam

OpenStudy (anonymous):

@kitteh♥ lol could you help please?

OpenStudy (cutiecomittee123):

Oh yikes, So are we just using geometric or arithmetic equations?

OpenStudy (cutiecomittee123):

If M was equal to one month then I think you could write their equations like so; Nita: y=5m+18 Tom: y=m+30

OpenStudy (cutiecomittee123):

Im not 100% sure but that is what I am guessing will help you.

OpenStudy (anonymous):

I got this equation:y=2x+30 y=4x+18

OpenStudy (anonymous):

So it would take 2 months